A gas cylinder containing cooking gas can whithstand a pressure of 14.9 atm . The pressure guaze of cylinder indicates 12 atm at 27 ℃. Due to sudden fire in the building tempr. Starts rising . The temperature at which cylinder will explode is :

Grade:12th pass

1 Answers

Arun
5 years ago

P1 = 12 atm

T1 = 300K

P2 =14.9 atm

Solution:

P1/T1 = P2/T2

Hence 12/300 = 14.9/T2

T2 = 372.5K

 

Regards

Arun (askIITians forum expert)